# Evaluating how LLM annotations represent diverse views on contentious topics

### v1

@@ 3.5 Data Processing and Sampling
To assess whether LLM models are more likely to align with majority demographic groups, we recoded gender, race, and education variables into binary categories: man/non-man, white/non-white, and non-college-educated/college-educated. This transformation allows us to analyze agreement patterns without focusing on the demographic predictors themselves, instead highlighting potential biases favoring dominant social groups.
@@
For each text, we calculated normalized label entropy by taking the set of human labels for a given text, calculating entropy, and dividing by the log number of labels. More specifically, Label $\text{Label Entropy}=\frac{-\sum_{x}^{n}p(x)\log p(x)}{log(n)}$, where $p(x)$ represents the probability of an annotator selecting a particular label and $n$ represents the total number of labels. We use label entropy as a measure of the disagreement among human annotators. Low label entropy indicates a high level of agreement across annotators, and high label entropy indicates a low level of agreement among annotators. Disagreement among annotators can be a result of task difficulty, where texts that do not neatly fit into one category or another have high levels of disagreement due to the difficulty of discernment. However, high label entropy can also indicate legitimate disagreement because of subjective views of the concept of interest.
Due to budget constraints, we used stratified samples of the Wikipedia Toxicity dataset. We stratified this sample based on label entropy, ensuring that our sample is balanced across texts of varying levels of disagreement among annotators.
3.6 LLM Annotations
@@
For this experiment, we employ GPT-4o Mini, Llama 3.3 70B Instruct, and Gemini 1.5 Flash. These models represent a diverse range of architectures, training processes, and openness, and are widely used in LLM research, enabling a broader analysis of model behavior across different language models researchers might use. To mitigate prompt-specific biases, we designed three distinct prompts for each dataset. These prompts can be found in the appendix. Each language model is provided with all three prompts, generating labels for each input text accordingly. [^1] For the LLM annotations, we operationalized “toxicity” as our annotation framework to bridge the methodological gap between the narrow “hate speech” criteria used by human coders and the broader “rude, hateful, aggressive, or unreasonable language” criteria applied to LLMs in the original NLPositionality study. Utilizing three language models and three prompts, each text receives a total of nine independent LLM-generated annotations, resulting in a total of 75,223 annotations across all four tasks. For each of the LLM annotations, we extract the label from the response using GPT-4o mini (e.g. for the toxicity task, extracting “yes” from “yes, the text is toxic.”) We then compare the LLM annotations with human annotations, specifically evaluating whether the models’ annotation aligns with the human annotations.
@@
To examine factors influencing agreement between LLM and human annotations, we use logistic regression models. Our dependent variable is model-human agreement (where agreement is 1 and disagreement is 0). For independent variables, we include the demographic attributes of the human annotators to assess potential biases in model alignment. We also control for label entropy–which captures disagreement among human annotators–and prompt to account for any prompt-specific or data-specific effects on model responses. For the main results, we perform a logistic regression model for each dataset and language model. Thus, agreement is predicted as a function of label entropy, prompt, gender, race (where available), and education. To account for the number of regressions we run, we apply Bonferroni correction for significance testing.
4 Results
Figure 1 shows the regression results for each data set and LLM. We show the probabilities of an LLM agreeing with each demographic group, where blue-colored points mean that the LLM agrees with the majority demographic group, red points mean that the LLM agrees with the minority demographic group, and the size of the point denotes the coefficient size.
In the NLPositionality dataset, the LLMs were less likely to agree with college-educated annotators, annotators who identify as men, and white annotators. Within the POPQUORN offensiveness dataset, the LLMs were less likely to agree with college-educated annotators and more likely to agree with white annotators. GPT-4o mini was less likely to agree with man-identifying annotators, but Llama and Gemini did not exhibit a statistically significant likelihood of agreement with annotators who identify as men. For the POPQUORN politeness dataset, the LLMs were more likely to agree with college-educated annotators, less likely to agree with annotators who identify as men, and more likely to agree with white annotators, though some of these were not statistically significant. Lastly, for Wikipedia, LLMs agreed with annotators who identify as men, and results for education were insignificant. Taken together, the LLM agreements with annotators are generally consistent within a dataset, but they are not consistent across datasets.
@@ 4 Results
The different results across datasets indicate that the LLMs do not exhibit a general tendency to agree with one demographic group over another. Notably, we found that within datasets, different LLMs are generally consistent regarding the demographic group they agree with, indicating that agreement with different demographic groups is dataset-specific, and may be a result of the underlying data, task, or annotator pool. Additionally, the range of probabilities is between 0.46 and 0.55, indicating that even at their most biased, LLMs are only about 5% more likely to agree with one demographic group over another. This suggests that LLMs may not systematically over-represent or under-represent the views of particular demographic groups to a large degree.
@@
To further examine the factors contributing to LLM agreement with annotators, we use logistic regression to predict agreement as a function of label entropy (our measure for annotator disagreement), LLM used, annotator gender, annotator race, annotator education, and LLM prompt. Figure 2 shows the odds ratios for each regression for each dataset. As the figure shows, the strongest predictor, by far, for LLM agreement is label entropy. LLM agreement is inversely related to label entropy; as label entropy increases, the LLMs are less likely to agree with annotators. This intuitively makes sense: if annotators largely disagree with each other, then there are by default fewer possible annotators for the LLM to agree with. Additionally, texts with high amounts of disagreement can be considered texts that are trickier or more difficult to label, hence causing disagreement among the annotators and reducing the likelihood that the LLM would agree with any particular annotator. To a smaller and less consistent degree, LLM choice plays a role in agreement predictions. For the NLPositionality and Politeness datasets, Llama and Gemini lead to lower agreement than GPT-4o mini. For the POPQUORN offensiveness dataset, Gemini was less likely to produce labels that agreed with annotators than GPT-4o mini, while Llama produced labels with higher rates of agreement with annotators. For Wikipedia, Llama was less likely to agree with annotators than GPT-4o mini. In summary, the model does generally matter, but there is not a model that exhibits agreement with annotators more consistently across datasets. Prompts chosen also matter, but not consistently. In some cases the prompt used does influence agreement, but the impact of prompts is lower than that of label entropy and LLM. [^2] We estimated a model without entropy to assess whether it was masking demographic effects. Removing entropy from the model did not meaningfully change results. Details are in the Appendix.
@@
Figure 2: Odds Ratios Across Datasets: The figure shows the odds ratios from logistic regressions for each dataset predicting LLM agreement as a function of label entropy, LLM, annotator gender, annotator race, annotator education, and prompt. The odds ratios are displayed with their 95% confidence intervals. The x-axis is the odds ratio, and the y-axis is the variable of interest. Points are colored for each dataset: NLPositionality (red), POPQUORN Offensiveness (green), POPQUORN Politeness (red), and Wikipedia comments (purple).
5 Discussion
@@
We compared three LLMs’ annotations on four tasks and did not find clear evidence of bias toward any particular gender, race, or educational attainment demographic group. In some cases, the LLMs were biased towards majority demographic groups, and in other cases the models were biased towards minority demographic groups. Moreover, in cases where there was a statistically significant probability of an LLM agreeing with one demographic group relative to another, the probability of agreement was not large; the probabilities of agreeing with any demographic group were between 0.46 and 0.55. We did, however, find that the specific LLM, prompt, and difficulty of the task were significantly predictive of LLM agreement with human annotators.
@@
The core concern motivating our paper is that known biases in LLMs (Kotek et al., 2023; 2024; Omar et al., 2025; Hastings, 2024; Thakur, 2023) may manifest in privileging the views of some demographic groups over others in data annotations, impacting the statistical inferences that use those annotations. These findings suggest that these concerns are unfounded, at least in the datasets we analyzed. While we observed some variation in LLM agreement with different demographic groups, these patterns were dataset-specific rather than indicative of a generalized bias in the models. This is contrary to what we might expect from existing research on LLM bias, which demonstrates that LLMs embody and perpetuate stereotypes in their linguistic outputs (Kotek et al., 2024). Instead, perhaps due to the close-ended nature of the task, we find minimal evidence of LLMs agreeing with one demographic group over another. The lack of consistency in demographic agreement across datasets implies that any observed differences are more likely a function of the underlying annotator pool and data characteristics rather than inherent biases in the LLMs themselves.
5.1 Implications for Researchers
@@
For researchers and practitioners considering LLMs for data annotation, our results indicate that while LLMs do not appear to introduce strong demographic biases, challenges remain in handling ambiguous or highly subjective labeling tasks. Literature on persona effects (Hu & Collier, 2024) suggests persona-specific prompting can effectively increase agreement for particular demographic groups. However, our results indicate that there are diminishing returns to persona-specific prompting because the difference between demographic groups and LLMs’ agreement is already low. These small gains in accuracy are likely not worth the potential costs of persona-specific stereotyping those prompts can elicit. In some cases, the benefits of using LLMs to annotate data may outweigh representational harms; for example, it may be better to use LLMs to annotate data than relying on human annotators because the annotation task itself can cause significant harm to annotators. For instance, exposing annotators to hate speech or toxic content carries risks for those annotators (Li et al., 2024; Schöpke-Gonzalez et al., 2024).
By far, the most significant predictor of LLM disagreement with annotators was label entropy, or disagreement among human annotators, indicating that LLMs and human annotators struggled with similar annotation tasks. Contrary to our expectations in this study, we do not find consistent or substantial bias towards any demographic group in particular. While there may be other reasons that researchers are hesitant to use LLMs to annotate texts (e.g. concerns about the environment, budget, or replicability of closed-source models), our results do not indicate that underrepresenting particular demographic groups is a substantial concern.
5.2 Limitations and Future Work
@@
We are only looking at four datasets and four tasks, and given the heterogeneity of our findings, we are not able to provide clear evidence regarding bias in large language models on subjective annotation tasks. Moreover, we only test three LLMs, so other LLMs may not behave this way or actually may be rather biased. We also only test datasets in English, and our results may not hold in other language contexts. However, the consistency of our results within dataset would suggest that bias comes from the underlying data generating process rather than model annotations themselves given the uniformity of results within dataset. Additionally, for the closed source models we tested (GPT-4o mini and Gemini), models may be updated, meaning results may not be consistent over time (while open-source models may also change over time, researchers could use version control to use older versions of models) (Munger, 2023). In recognition of the potential harms of LLMs, LLM producers also use red-teaming and safety alignment to modify model outputs, which may also explain some of the results in this paper. However, due to lack of transparency from these producers, we cannot provide clear evidence one way or another. Future work should explore whether these findings hold across additional datasets and tasks, as well as investigate strategies for improving LLM performance in cases of high annotator disagreement. Additionally, while LLMs may not systematically favor particular demographic groups, their responses are still influenced by model choice and prompting strategy, underscoring the need for careful prompt engineering and validation when using these models for annotation in research and applied settings.
6 Conclusion
@@
Our study evaluates whether large language models (LLMs) exhibit systematic bias in agreement with annotators from different demographic groups. Across four annotation tasks using four datasets and three LLMs, we found that LLM agreement with annotators was not primarily driven by demographic differences. Instead, the strongest predictor of LLM agreement was label entropy, or the level of disagreement among human annotators on the original annotation task. This suggests that rather than consistently favoring particular demographic groups, LLMs struggle most with annotating texts that human annotators themselves find ambiguous or contentious. Additionally, we found that the choice of LLM and the specific prompt used to elicit labels also influenced agreement rates, but no single model consistently outperformed the others across all tasks. These results suggest that LLMs can be used to label social science data for research without substantial representational harms.
Appendix
Appendix A Robustness Checks
@@
To ensure that label entropy is not masking demographic variation, we run the same logistic regressions as specified in Figure 1, but we leave label entropy out of the controls. The results remain largely the same as our main specification, alieviating concerns that controlling for label entropy is masking true effects. The results can be found in Figure LABEL:fig:regdemographicsappendix.
@@
Appendix B Classification Prompts
We used three semantically-equivalent classification prompts with each LLM.
B.1 NLPositionality Toxicity Data
Prompt 1:
I want you to perform a data annotation task. Your task is to carefully read the text provided below and determine whether the text is toxic or not. In your output, I want you to respond with yes if the text is toxic, otherwise respond with no.
Prompt 2:
Perform data annotation by determining if the text is toxic or not. Only respond with yes or no.
Prompt 3:
Carefully review the text below and assess whether it is toxic. If the text is toxic, respond with yes; if not, respond with no.
B.2 POPQUORN Offensiveness Data
Prompt 1:
I want you to perform an annotation task. Your task is to carefully read the text provided below and determine how offensive the text is on a scale of 1-5, where ”1” is ”not offensive at all” and 5 is ”very offensive.”
Prompt 2:
Complete an annotation task by carefully reading the text below and rating its offensiveness on a scale of 1 to 5, where ”1” means ”not offensive at all” and ”5” indicates ”very offensive.”
Prompt 3:
Rate the offensiveness of the text below from 1 (not offensive) to 5 (very offensive).
B.3 POPQUORN Politeness Data
Prompt 1:
I want you to perform an annotation task. Your task is to carefully read the email provided below and determine how polite the email is on a scale of 1-5, where ”1” is ”not polite at all” and 5 is ”very polite.”
Prompt 2:
Complete an annotation task by carefully reading the text below and rating its politeness on a scale of 1 to 5, where ”1” means ”not polite at all” and ”5” indicates ”very polite.”
Prompt 3:
Rate the politeness of the text below from 1 (not polite) to 5 (very polite).
B.4 Wikipedia Toxicity Data
Prompt 1:
I want you to perform a data annotation task. Your task is to carefully read the text provided below and determine whether the text is toxic or not. In your output, I want you to respond with yes if the text is toxic, otherwise respond with no.
Prompt 2:
Perform data annotation by determining if the text is toxic or not. Only respond with yes or no.
Prompt 3:
Carefully review the text below and assess whether it is toxic. If the text is toxic, respond with yes; if not, respond with no.
Appendix C Extraction Prompts
Previous studies have shown that prompting LLMs to be concise can reduce accuracy in response (Deng et al., 2024). For the POPQUORN Offensiveness and Politeness datasets, we used an extraction step to obtain the LLMs’ ratings (from 1 to 5) from their responses to the classification prompts. Specifically, we used the following extraction prompt for the Offensiveness dataset:
Based on your response above, what is the numeric rating on the offensiveness scale? Only return a number with no other words, punctuation, or characters. If no number is provided, respond with ”N/A”.

### v2

@@ III-E Data Processing and Sampling
To assess whether LLM models are more likely to align with socially dominant demographic groups, we recoded gender, race, and education variables into binary categories: man/non-man, white/non-white, and college-educated/non-college-educated. This transformation was undertaken to make the datasets comparable, as each dataset had a different breakdown of each demographic. Such a transformation also allowed us to analyze agreement patterns without focusing on the demographic predictors themselves, instead highlighting potential biases favoring dominant social groups.
For each text, we calculated the normalized label entropy by taking the set of human labels for a given text, calculating entropy, and dividing by the log number of labels. More specifically,
Label $\text{Label Entropy}=\frac{-\sum_{x}^{n}p(x)\log p(x)}{\log(n)}$
@@
where $p(x)$ represents the probability of an annotator selecting a particular label and $n$ represents the total number of labels. We used label entropy as a measure of disagreement among human annotators. A low label entropy indicated a high level of agreement across annotators; a high label entropy indicated a low level of agreement among annotators. Disagreement among annotators can be a result of task difficulty, where texts that do not neatly fit into one category or another have high levels of disagreement due to the difficulty of discernment. However, high label entropy can also indicate legitimate disagreement because of subjective views of the concept of interest.
III-F LLM Annotations
@@
For the experiments, we employed GPT-4o Mini, Llama 3.3 70B Instruct, and Gemini 1.5 Flash. These models represent a diverse range of architectures, training processes, and openness, and are widely used in LLM research, enabling a broader analysis of model behavior across different LLMs researchers commonly use. To mitigate prompt-specific biases, we designed three distinct but semantically equivalent prompts for each dataset. These prompts can be found in the Appendix. Each language model was prompted using each semantically equivalent prompt, generating labels for each input text accordingly. [^1] For the LLM annotations, we operationalized “toxicity” as our annotation framework to bridge the methodological gap between the narrow “hate speech” criteria used by human coders and the broader “rude, hateful, aggressive, or unreasonable language” criteria applied to LLMs in the original NLPositionality study. Utilizing three language models and three prompts, each text received a total of nine independent LLM-generated annotations, resulting in a total of 75,223 annotations across all four tasks. For each of the LLM annotations, we extracted the label from the response using GPT-4o mini (e.g. for the toxicity task, extracting “yes” from “yes, the text is toxic.”) We then compared the LLM annotations with human annotations, specifically evaluating whether the models’ annotations aligned with the human annotations. All experiments were run using OpenAI’s API for GPT-4o mini, DeepInfra for Llama 3.3 70B Instruct, and the Google Gemini Developer API for Gemini 1.5 Flash. We used default hyperparameters (temperature and top-$p$) for all models.
@@
To examine factors influencing agreement between LLM and human annotations, we used logistic regression models. Specifically, our dependent variable was model-human agreement, where agreement is 1 and disagreement is 0, and our independent variables were the demographic attributes of the human annotators. We also controlled for label entropy—which captures disagreement among human annotators—and prompt variation to account for any prompt-specific or data-specific effects on model responses. In other words, our logistic regression was the following.
Agreement $=\beta_{0}+\beta_{entropy}\text{Label Entropy}$
[…]
To account for the number of regressions we ran, we applied Bonferroni correction for significance testing. For the main results, we performed a logistic regression model for each dataset and language model.
IV Results
Figure 1 shows the logistic regression results for each data set and LLM. We present both the logistic regression coefficients and their corresponding probabilities of LLM agreement with each demographic group; the latter are derived from the regression coefficients and provide more intuitive interpretation of the results. Colored points indicate statistically significant coefficients with $\alpha<0.05$ (after Bonferroni corrections); blank points indicate non-significant coefficients. A blue-colored point indicates that the LLM’s generated responses agreed with the socially dominant demographic group (i.e., the reference category); a red point means that the LLM’s generated responses agreed with the less socially dominant demographic group. The size of the point denotes the coefficient’s magnitude.
@@ IV Results
In the NLPositionality dataset, all LLMs were less likely to agree with college-educated annotators, annotators who identify as men, and white annotators. Within the POPQUORN offensiveness dataset, all LLMs were less likely to agree with college-educated annotators and more likely to agree with white annotators. GPT-4o mini was less likely to agree with annotators identifying as men, but Llama and Gemini did not exhibit a statistically significant likelihood of agreement with annotators who identify as men. For the POPQUORN politeness dataset, the LLMs were more likely to agree with college-educated annotators, less likely to agree with annotators who identify as men, and more likely to agree with white annotators, though some of these were not statistically significant. Lastly, for Wikipedia, the LLMs agreed with annotators who identify as men, and results for education were insignificant.
@@
Taken together, the way LLMs’ generated responses agreed with annotators with specific demographics were generally consistent within a dataset, but were not model-specific. In other words, we found that within datasets, different LLMs were generally consistent regarding the demographic group they agreed with, indicating that agreement with different demographic groups was dataset-specific. This consistency suggests that the task or domain drives the bias more than the choice of LLM. We also found that the range of probabilities was between 0.46 and 0.55, indicating that even at their most biased, LLMs were only about 5% more likely to agree with one demographic group over another. These different results across datasets suggest that LLMs do not exhibit a general tendency to systematically over-represent or under-represent the views of particular demographic groups, but rather may do so depending on the specific annotation task at hand.
@@
To further examine the factors contributing to LLM agreement with annotators, we turned to another logistic regression that used the same set of predictors as before along a categorical variable that coded the LLM used, allowing us to understand the influence of LLM choice on agreement. Figure 2 shows the odds ratios for each predictor for each dataset. As the figure shows, the strongest predictor for LLM agreement, by far, was label entropy. LLM agreement was inversely related to label entropy—as label entropy increased, the LLMs were less likely to agree with annotators. This intuitively makes sense: if annotators largely disagree with each other, then there are by default fewer possible annotators for the LLM to agree with. Additionally, texts with higher levels of disagreement can be considered texts trickier or more difficult to label, hence causing disagreement among the annotators and reducing the likelihood that the LLM would agree with any particular annotator.
@@
Figure 2: Odds Ratios Across Datasets: The figure shows the odds ratios from logistic regressions for each dataset predicting LLM agreement as a function of label entropy, LLM, annotator gender, annotator race, annotator education, and prompt. The odds ratios are displayed with their 95% confidence intervals. The x-axis is the odds ratio, and the y-axis is the variable of interest. Points are colored for each dataset: NLPositionality (red), POPQUORN Offensiveness (green), POPQUORN Politeness (blue), and Wikipedia comments (purple).
To a smaller and less consistent degree, the LLM choice played a role in the agreement predictions. For the NLPositionality and Politeness datasets, Llama and Gemini led to lower agreement than GPT-4o mini. For the POPQUORN offensiveness dataset, Gemini was less likely to produce labels that agreed with annotators than GPT-4o mini, while Llama produced labels with higher rates of agreement with annotators. For Wikipedia, Llama was less likely to agree with annotators than GPT-4o mini. In short, the choice of the model did generally matter, but, consistent with our findings in our previous logistic regressions, there was no model that demonstrated agreement with annotators more consistently across all datasets.
The semantically equivalent prompt used also mattered, but not consistently. Overall, the impact of the chosen prompt was lower than that of label entropy and the LLM used.
In short, label entropy was by far the most predictive factor of agreement between LLMs and human annotators. While there were significant odds ratios with model choice and prompt choice, the effect sizes were by far smaller than the effect sizes of label entropy. We note that we estimated a model without entropy to assess whether it was masking demographic effects. Removing entropy from the model did not meaningfully change results. Details are in the Appendix.
V Discussion
@@
We compared three LLMs’ annotations on four tasks and did not find clear evidence of systematic bias toward any particular gender, race, or educational attainment demographic group. Instead, we found that the bias was dataset-specific. All LLMs tended to be biased in the same direction on each dataset. This bias was towards socially dominant demographic groups on some datasets, and in other datasets, the models were biased towards less socially dominant demographic groups. We also found that in cases where there was a statistically significant probability of an LLM agreeing with one demographic group relative to another, the probability of agreement was relatively small: the probabilities of agreeing with any demographic group were between 0.46 and 0.55. We found, instead, that the difficulty of the task was a much more significant predictor of the LLM agreement with human annotators.
@@
The core concern motivating our paper is that known biases in LLMs [14, 15, 13, 22, 12] may manifest themselves in privileging the views of some demographic groups over others in data annotations, impacting the statistical inferences that use those annotations. Our findings suggest that these concerns are present, as demonstrated by the statistically significant coefficients in our logistic regressions, but they manifest in a dataset-specific manner. The patterns are not indicative of a generalized, systematic bias in the models towards a particular set of demographic groups. These findings highlight the nuance of LLM bias. It can be contrary to what we might expect from existing research on LLM bias, which demonstrates that LLMs embody and perpetuate stereotypes in their linguistic outputs [15]. The lack of consistency in demographic agreement across datasets implies that any observed differences are more likely a function of the data characteristics and how it interacts with the LLM’s pretrained embedded information rather than inherent and systematic biases embedded in the LLMs that manifest in all subjective annotation tasks. This finding may also be related to recent findings that representations in AI models are converging, but explorations of this connection are outside the scope of this paper (see, e.g., [45]).
V-A Implications for Researchers and Practitioners
For researchers and practitioners considering using generative LLMs for data annotation, our results indicate that while LLMs do not appear to introduce strong systematic demographic biases, challenges remain in handling ambiguous or highly subjective labeling tasks. We highlight four implications for researchers.
Fairness evaluations must be contextual
A specific LLM can appear pro-minority in one setting and pro-majority in another. The specific demographic group with which the LLM most strongly agrees, whether that is an annotator of a specific race, gender, demographic, or some combination of the three, may also vary by data set. We suggest that practitioners use a demographically balanced evaluation dataset to identify potential biases of the generative LLM used.
Model choice alone will not solve bias
As our findings indicate, LLMs tend to be biased in the same directions within the same dataset. Turning to a different LLM with a different architecture or training strategy may not effectively mitigate biases.
Persona-specific prompting may yield diminishing returns
@@ V-A Implications for Researchers and Practitioners
The literature on persona effects [26] suggests persona-specific prompting can effectively increase agreement for particular demographic groups. However, our results indicate that there are diminishing returns to persona-specific prompting because the difference between demographic groups and LLMs’ agreement is already relatively low. These small gains in accuracy are likely not worth the potential costs of persona-specific stereotyping those prompts can elicit. In some cases, the benefits of using LLMs to annotate data may outweigh representational harms. For example, it may be better to use LLMs to annotate data than relying on human annotators because the annotation task itself can cause significant harm to annotators. Exposing annotators to hate speech or toxic content carries risks for those annotators [33, 46].
Effect sizes of bias are often significant but small, and item difficulty is usually the much stronger predictor of human disagreement with LLMs
It is important to balance the evaluation of biases with the fact that some texts, especially in subjective annotation tasks, are simply more difficult to label. In our results, the most significant predictor of LLM disagreement with annotators was label entropy, or disagreement among human annotators, indicating that LLMs and human annotators struggled with similar annotation tasks. It is critical to explicitly model item difficulty when examining potential demographic biases with the LLM’s generated labels.
V-B Limitations and Future Work
@@
We are only looking at four datasets, and given the heterogeneity of our findings, we are unable to provide clear evidence regarding systematic, consistent biases in some of the most popular LLMs on subjective annotation tasks. Moreover, we only test three LLMs, so other LLMs may not behave this way or actually may be rather biased. We also only test datasets in English, and our results may not hold in other language contexts. Additionally, for the closed source models we tested (GPT-4o mini and Gemini), models may be updated, meaning results may not be consistent over time. While open weight models may also change over time, researchers could use version control to use older versions of models [47]. In recognition of the potential harms of LLMs, LLM developers also use red-teaming and safety alignment techniques to regulate model outputs, which may also explain some of the results in this paper. However, due to lack of transparency from these developers about how these models were trained and what data was used, we cannot provide clear evidence on the effects of red-teaming and safety alignment on our results. Future work should explore whether these findings hold across additional datasets and tasks, as well as investigate strategies for improving LLM performance in cases of high annotator disagreement. Additionally, while LLMs may not systematically favor particular demographic groups, their responses are still influenced by model choice and prompting strategy, underscoring the need for careful prompt engineering and validation when using these models for annotation in research and applied settings.
VI Conclusion
@@
Our study evaluated whether large language models (LLMs) exhibit systematic bias in agreement with annotators from different demographic groups. Across four annotation tasks using four datasets and three LLMs, we found that LLM agreement with annotators was not systematically driven by specific demographic differences. Instead, LLMs tended to exhibit biases in the same directions within each dataset; in other words, the biases exhibited are not model-specific, but dataset-specific. Moreover, by far the strongest predictor of LLM agreement was label entropy, or the level of disagreement among human annotators on the original annotation task. This suggests that rather than consistently favoring particular demographic groups, LLMs struggle most with annotating texts that human annotators themselves find ambiguous or contentious. Additionally, we found that the choice of LLM and the specific prompt used to elicit labels also influenced agreement rates, but no single model or prompt consistently outperformed the others across all tasks. Practically, our results suggest that fairness evaluations must be contextual, model choice alone will not solve bias, and item difficulty must be integrated into bias assessments.
Appendix A Appendix
A-A Robustness Checks
@@
To ensure that label entropy is not masking demographic variation, we run the same logistic regressions as specified in Figure 1, but we leave label entropy out of the controls. The results remain largely the same as our main specification, alleviating concerns that controlling for label entropy is masking true effects. The results can be found in Figure 3.
@@
A-B Classification Prompts
We used three semantically equivalent classification prompts with each LLM.
NLPositionality Toxicity Data
  • Prompt 1: I want you to perform a data annotation task. Your task is to carefully read the text provided below and determine whether the text is toxic or not. In your output, I want you to respond with yes if the text is toxic, otherwise respond with no.
  • Prompt 2: Perform data annotation by determining if the text is toxic or not. Only respond with yes or no.
  • Prompt 3: Carefully review the text below and assess whether it is toxic. If the text is toxic, respond with yes; if not, respond with no.
POPQUORN Offensiveness Data
  • Prompt 1: I want you to perform an annotation task. Your task is to carefully read the text provided below and determine how offensive the text is on a scale of 1-5, where “1” is “not offensive at all” and 5 is “very offensive.”
  • Prompt 2: Complete an annotation task by carefully reading the text below and rating its offensiveness on a scale of 1 to 5, where “1” means “not offensive at all” and “5” indicates “very offensive.”
  • Prompt 3: Rate the offensiveness of the text below from 1 (not offensive) to 5 (very offensive).
POPQUORN Politeness Data
  • Prompt 1: I want you to perform an annotation task. Your task is to carefully read the email provided below and determine how polite the email is on a scale of 1-5, where “1” is “not polite at all” and 5 is “very polite.”
  • Prompt 2: Complete an annotation task by carefully reading the text below and rating its politeness on a scale of 1 to 5, where “1” means “not polite at all” and “5” indicates “very polite.”
  • Prompt 3: Rate the politeness of the text below from 1 (not polite) to 5 (very polite).
Wikipedia Toxicity Data
  • Prompt 1: I want you to perform a data annotation task. Your task is to carefully read the text provided below and determine whether the text is toxic or not. In your output, I want you to respond with yes if the text is toxic, otherwise respond with no.
  • Prompt 2: Perform data annotation by determining if the text is toxic or not. Only respond with yes or no.
  • Prompt 3: Carefully review the text below and assess whether it is toxic. If the text is toxic, respond with yes; if not, respond with no.
A-C Extraction Prompts
Previous studies have shown that prompting LLMs to be concise or prompting LLMs into a forced choice can reduce accuracy in response [48, 48]. For the POPQUORN Offensiveness and Politeness datasets, we used an extraction step to obtain the LLMs’ ratings (from 1 to 5) from their responses to the classification prompts. Specifically, we used the following extraction prompt for the Offensiveness dataset:
Based on your response above, what is the numeric rating on the offensiveness scale? Only return a number with no other words, punctuation, or characters. If no number is provided, respond with “N/A”.